3 tables and 2 chairs cost Rs. 1900 and 2 tables and 4 chairs cost Rs. 1800. Find the cost of table and a chair.
Answer by Cromlix(4381)

Hi there,
Make tables = x
and chairs = y
Make two equations:
3x + 2y = 1900.....(1)
2x + 4y = 1800.....(2)
Multiply equation (1) by 2
6x + 4y = 3800.....(1)
2x + 4y = 1800
Subtract Equation (2) from (1)
4x = 2000
x = 500
Substitute x = 500 into
Equation (2)
2x + 4y = 1800
2(500) + 4y = 1800
1000 + 4y = 1800
4y = 1800 - 1000
4y = 800
y = 200
Therefore:
A Table costs Rs 500
A chair costs Rs 200
Hope this helps:-)